### v1

@@ 3.3 The $\bar{\mu}$-distance as a metric
By the definition as set out in Section 2 the symmetric difference between two multisets is empty, if, and only if, they are the same multiset. Furthermore, because the $\mu$-representation of a network is well-defined, if two networks are isomorphic their $\mu$-representations are equal. If however, two networks have equal $\mu$-representation, this does not necessarily mean they are isomorphic, see the examples in Figure 1 and Figure 7. Theorem 1 shows that given two semi-binary stack-free networks with equal modified $\mu$-representations, if one of them is orchard, then they are isomorphic. Let us define the $\bar{\mu}$-distance on networks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ by taking the cardinality of the symmetric difference of modified $\mu$-representations, i.e., $d_{\bar{\mu}}(\mathcal{N}_{1},\mathcal{N}_{2})=|\bar{\mu}(\mathcal{N}_{1})\triangle\bar{\mu}(\mathcal{N}_{2})|$. By Theorem 1, this is a metric on the class of semi-binary stack-free orchard networks.
@@ 3.4 Non-binary stack-free orchard networks
In this section we will discuss whether our encoding results for the modified $\mu$-representation extend to non-binary orchard networks. First, we suppose Observation 1 regarding the parents of leaf nodes, the multiplicity of the $\mu$-vectors of leaf nodes, and the $\mu$-vector of the tree-node parent of a leaf node, holds for non-binary stack-free networks without any further modification. Similarly, we suggest Lemma 1 holds for non-binary stack-free networks and therefore cherries are uniquely determined by the $\mu$-representation for non-binary stack-free networks.
However, Lemma 2 does not have an obvious equivalent for non-binary networks. When considering whether the leaf pair $(b,a)$ is a reticulated cherry, we can no longer require the existence of $\mu(p_{a})=\mu(a)+\mu(b)$, because the parent of $a$ may have more children than just $a$ and $p_{b}$. By Observation 1 we can find the $\mu$-vector of the parent of $a$, but if it is not equal to $\mu(a)+\mu(b)$, then it is impossible to determine whether there are any other nodes on the path $p_{a}\rightsquigarrow p_{b}$. Figure 8 displays two non-binary stack-free orchard networks with the same $\bar{\mu}$-representation, which are not isomorphic. Because the parent $p_{a}$ of $a$ has three children and its $\mu$-vector $\mu(p_{a})$ is equal to the sums of the $\mu$-vectors of 2 different sets of 3 nodes ($1111=1000+0100+0011=1000+0110+0001$) of which all those that differ belong to reticulations, it is impossible to tell which set belongs to the children of $p_{a}$ and therefore whether $(b,a)$ is a reticulated cherry or not. As a consequence of the example given in Figure 8 we obtain the following:
@@ 4.1 Determining the in-degrees of reticulations
In [CRV08], Cardona et al. showed that the $\mu$-representation serves as an encoding for non-binary tree-child phylogenetic networks. However, we have shown that for more general networks, even if they are stack-free semi-binary, two non-isomorphic networks may have the same $\mu$-representation. See Figure 1 for an example. We have also shown that two semi-binary stack-free networks which have the same $\mu$-representation are isomorphic, as long as one is orchard and their nodes have the same in-degrees (Theorem 1). Therefore, if we can find a subclass of semi-binary stack-free networks, for which equivalent $\mu$-representations imply the same in-degrees, then we can show that equivalent $\mu$-representations imply isomorphism as long as one of the networks is orchard. This would give us an encoding result for this subclass, which would mean that the symmetric difference of the $\mu$-representations can be used as a metric. In this and the next sections we will investigate the relation between $\mu$-vectors and in-degrees and show that there is a subclass for which the in-degrees are uniquely determined by the $\mu$-representation.
@@ 4.1.1 An equation relating the in-degrees of reticulations and $\mu$-vectors
Let $\mathcal{N}=(V,E)$ be a non-binary X-DAG.
@@
Let $\mathcal{N}$ be an $X$-DAG and let $a$ be an element of X. Let $R$ be the set of reticulations in $V$. Then,
$\mu_{a}(\rho)=\sum_{r_{i}\in R}(\delta^{-}(r_{i})-1)\mu_{a}(r_{i})+1.$
Note that $\mu_{a}(r_{i})=0$ if and only if $r_{i}$ is not an ancestor of $a$, which means that the contributions to the sum of Proposition 1 come from reticulation ancestors of $a$.
@@
Let us consider an $X$-DAG with $k=0$, i.e., a tree. If the hybridization number is zero then the graph contains no reticulations and therefore $R$ is empty. This also means there is a unique tree-path from the root to each leaf and thus $\mu_{a}(\rho)=1$. This shows the equation holds for each leaf of this graph.
Now suppose the equation holds for each leaf of any $X$-DAG with hybridization number lower than $k$. Let $\mathcal{N}$ be an $X$-DAG with reticulation set $R$ such that $h(\mathcal{N})=k$. Without loss of generality let $r_{j}$ be a highest reticulation in $\mathcal{N}$, which means $r_{j}$ has no reticulation ancestors. We can decrease the in-degree of $r_{j}$ by deleting an incoming edge $ur_{j}$. If $u$ is an elementary node we delete all edges on the maximal elementary path which visits $u$, as well as the nodes which become isolated by doing so. When we delete any of the incoming edges $ur_{j}$ in this way the resulting network $\mathcal{N}^{\prime}$ is still an $X$-DAG. Furthermore, the in-degrees of any other reticulations in the network have not decreased, which means $\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})=\delta^{-}_{\mathcal{N}}(r_{i})$ for any $i\neq j$. Moreover, the number of paths from any reticulation $r$ to any leaf $a$ has not changed and thus $\mu^{\prime}_{a}(r)=\mu_{a}(r)$. Finally, the hybridization number of $\mathcal{N}^{\prime}$ is equal to $k-1$ and by the induction hypothesis we have,
$\mu_{a}^{\prime}(\rho)=\sum_{r_{i}\in R^{\prime}}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu_{a}(r_{i})+1.$
@@
$\mu_{a}^{\prime}(\rho)$ $=\sum_{r_{i}\in R^{\prime}}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu_{a}(r_{i})+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu_{a}(r_{i})+(\delta^{-}_{\mathcal{N}}(r_{j})-2)\mu_{a}(r_{j})+1$
@@
$\mu_{a}^{\prime}(\rho)$ $=\sum_{r_{i}\in R^{\prime}}((\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu_{a}(r_{i})+1$
$=\sum_{r_{i}\in R}((\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu_{a}(r_{i})+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu_{a}(r_{i})+(\delta^{-}_{\mathcal{N}^{\prime}}(r_{j})-1)\mu_{a}(r_{j})+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu_{a}(r_{i})+(\delta^{-}_{\mathcal{N}}(r_{j})-2)\mu_{a}(r_{j})+1.$
@@
When we add the edge $ur_{j}$ (or the maximal elementary path which visits $u$ and ends in $r_{j}$) back to the network, we generate $\mathcal{N}$ from $\mathcal{N}^{\prime}$. In doing so we increase $\mu^{\prime}_{a}(\rho)$ by $\mu_{a}(r_{j})$ for any leaf $a$. Indeed, $r_{j}$ is chosen to be a highest reticulation in $\mathcal{N}$, and so there is a single path from the root to $r_{j}$ which uses the edge $(u,r_{j})$. By definition, there are $\mu_{a}(r_{j})$ paths from $r_{j}$ to any leaf $a$. This means that, for any leaf $a$
$\mu_{a}(\rho)$ $=\mu_{a}^{\prime}(\rho)+\mu_{a}(r_{j})$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu_{a}(r_{i})+(\delta^{-}(r_{j})_{\mathcal{N}}-2)\mu_{a}(r_{j})+1+\mu_{a}(r_{j})$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu_{a}(r_{i})+(\delta^{-}(r_{j})_{\mathcal{N}}-1)\mu_{a}(r_{j})+1$
$=\sum_{r_{i}\in R}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu_{a}(r_{i})+1,$
@@
Note that this theorem holds for non-binary $X$-DAGs, as we never assumed a limit on the in-degree of reticulations or the out-degree of tree-nodes. Proposition 1 provides us with a system of linear equations that govern the in-degrees of reticulations as a function of $\mu$-vectors. We know this system of equations must have a solution as long as it belongs to a valid phylogenetic network. However, there are no guarantees yet that it has a unique solution. A network may contain more reticulations than there are linearly independent equations in the system. The network displayed in Figure 9 is semi-binary stack-free orchard, yet the system of equations generated by applying Proposition 1 to each leaf does not have a unique solution. Note that leaf $c$ has no reticulation ancestors and therefore the equation for leaf $c$ ($\mu_{c}(\rho)=1$) does not contribute. This means that there is one equation for leaf $a$ and one for leaf $b$, but three variables, the in-degrees of reticulations $r_{1},r_{2}$ and $r_{3}$. Note that the lower bound for the in-degree of reticulations is 2, and therefore, there are only a finite number of solutions. In this case there are two solutions to the system of equations and Figure 9 shows $(\delta^{-}(r_{1}),\delta^{-}(r_{2}),\delta^{-}(r_{3}))=(3,3,2)$ is the one belonging to this network.
Figure 9: A semi-binary stack-free orchard network for which the system of equations generated by applying Proposition 1 to each leaf does not have a unique solution. However, the solution $(\delta^{-}(r_{1}),\delta^{-}(r_{2}),\delta^{-}(r_{3}))=(3,3,2)$ is the unique one that belongs to this network.
@@ 4.1.2 In-degrees of stable reticulations
If we let $\mathcal{N}$ be a reticulation-visible X-DAG with reticulation set $R$ the following is true. For $r\in R$, with leaf $a$ below $r$ such that all paths from the root to $a$ visit $r$, and $v$ any ancestor of $r$,
$\mu_{a}(v)=P_{vr}\mu_{a}(r),$ (1)
where $P_{vr}$ is the number of paths from $v$ to $r$. This is true, as every path from $v$ to $a$ must be the product of a path from $v$ to $r$ and a path from $r$ to $a$. With this we gain the following lemma.
@@
$\frac{\mu_{a}(\rho)}{\mu_{a}(r_{\ell})}$ $=\sum_{r\in A}(\delta^{-}(r)-1)\frac{\mu_{a}(r)}{\mu_{a}(r_{\ell})}+1$ (2)
@@
Let us generate $\mathcal{N}^{\prime}$ from $\mathcal{N}$ by first attaching a new leaf $a^{\prime}$ to $r_{\ell}$, by adding the edge $(r_{\ell},a^{\prime})$, and then adjusting the $\mu$-representation by adding a column for $a^{\prime}$, such that $\mu^{\prime}(v)=\mu(v)\oplus\mu_{a^{\prime}}(v)$. Note that this network is now an $X^{\prime}$-DAG, where $X^{\prime}=X\cup\{a^{\prime}\}$. Then by Proposition 1 we have:
@@
$\mu^{\prime}_{a^{\prime}}(v)=P_{vr_{\ell}}=\frac{\mu_{a}(v)}{\mu_{a}(r_{\ell})}.$ (4)
@@
$\frac{\mu_{a}(\rho)}{\mu_{a}(r_{\ell})}=\sum_{r\in A}(\delta^{-}(r)-1)\frac{\mu_{a}(r)}{\mu_{a}(r_{\ell})}+1.$
For the following lemmas, let $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ be two reticulation-visible X-DAG’s with $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$. Let both networks have the same reticulation set $R$ and the same leaf set $X$ Finally, assume that for each reticulation $r\in R$, there exists a leaf $a\in X$ such that $r$ is stable with respect to $a$ in both networks.
@@
For any reticulation $r\in H$ the set $A$ in Lemma 4 contains only $r$ and therefore the in-degree of $r$ is given directly by Equation 2, applied to the leaf for which the reticulation is stable in both networks. ∎
@@
Let $H$ be the set of highest reticulations in $R$, which is the same set for $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, by Lemma 5. By Lemma 6 the in-degrees of the reticulations in $H$ are the same in both networks. Let $S$ be the set of second highest reticulations in $R$, such that each reticulation $r\in S$ only has ancestors in $H$. By applying Lemma 7 the in-degrees of the reticulations in $S$ are also the same in both networks. Now, in the same way, the in-degrees of the third highest reticulations are equal and so on and so forth. This process must terminate as we only consider finite graphs. Thus, by repeated application of Lemma 7, we see that the in-degrees of all reticulations in $R$ are the same in both networks. ∎

### v2

@@ 3.3 The $\bar{\mu}$-distance as a metric
By the definition as set out in Section 2 the symmetric difference between two multisets is empty, if, and only if, they are the same multiset. Furthermore, because the $\mu$-representation of a network is well-defined, if two networks are isomorphic then their $\mu$-representations are equal. If however, two networks have equal $\mu$-representation, this does not necessarily mean they are isomorphic, see the examples in Figure 1 and Figure 7. Theorem 1 shows that given two semi-binary stack-free networks with equal modified $\mu$-representations, if one of them is orchard, then they are isomorphic. Let us define the $\bar{\mu}$-distance on networks $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ by taking the cardinality of the symmetric difference of modified $\mu$-representations, i.e., $d_{\bar{\mu}}(\mathcal{N}_{1},\mathcal{N}_{2})=|\bar{\mu}(\mathcal{N}_{1})\triangle\bar{\mu}(\mathcal{N}_{2})|$. By Theorem 1, this is a metric on the class of semi-binary stack-free orchard networks.
@@ 3.4 Non-binary stack-free orchard networks
In this section we will discuss whether our encoding results for the modified $\mu$-representation extend to non-binary orchard networks. We claim that 1 regarding the parents of leaf nodes, the multiplicity of the $\mu$-vectors of leaf nodes, and the $\mu$-vector of the tree-node parent of a leaf node, holds for non-binary stack-free networks without any further modification. Similarly, we claim 1 holds for non-binary stack-free networks and therefore cherries are uniquely determined by the $\mu$-representation for non-binary stack-free networks.
However, 2 does not have an obvious equivalent for non-binary networks. When considering whether the leaf pair $(b,a)$ is a reticulated cherry, we can no longer require the existence of $\mu(p_{a})=\mu(a)+\mu(b)$, because the parent of $a$ may have more children than just $a$ and $p_{b}$. By 1 we can find the $\mu$-vector of the parent of $a$, but if it is not equal to $\mu(a)+\mu(b)$, then it is impossible to determine whether there are any other nodes on the path $p_{a}\rightsquigarrow p_{b}$. Figure 8 displays two non-binary stack-free orchard networks with the same $\bar{\mu}$-representation, which are not isomorphic. Because the parent $p_{a}$ of $a$ has three children and its $\mu$-vector $\mu(p_{a})$ is equal to the sums of the $\mu$-vectors of 2 different sets of 3 nodes ($1111=1000+0100+0011=1000+0110+0001$) of which all those that differ belong to reticulations, it is impossible to tell which set belongs to the children of $p_{a}$ and therefore whether $(b,a)$ is a reticulated cherry or not. As a consequence of the counterexample given in Figure 8 we obtain the following:
@@ 4 Encoding by $\mu$-representations
In [CRV08], Cardona et al. showed that the $\mu$-representation serves as an encoding for non-binary tree-child phylogenetic networks. We have shown that for more general networks, even if they are stack-free semi-binary, two non-isomorphic networks may have the same $\mu$-representation. See Figure 1 for an example. We have also shown that for two semi-binary stack-free networks $\mathcal{N},\mathcal{N}^{\prime}$ with $\bar{\mu}(\mathcal{N})=\bar{\mu}(\mathcal{N}^{\prime})$ if either of them is orchard then they are isomorphic (Theorem 1). Therefore, if we can find a subclass where $\mu(\mathcal{N})=\mu(\mathcal{N}^{\prime})$ implies that the nodes in $\mathcal{N}$ and $\mathcal{N}^{\prime}$ have the same in-degrees then we can show that equivalent $\mu$-representations imply isomorphism as long as one of the networks is orchard. This would give us an encoding result for this subclass. In this section we set out to show the following main result.
Theorem (4).
We want to show that there exists a subclass of phylogenetic networks, which we call strongly reticulation-visible, that can be encoded by the $\mu$-representation without modification. To do so, we first give a set of equations which relate the in-degrees of the nodes in a phylogenetic network to the $\mu$-representation (Proposition 1). We then show that for reticulation-visible networks, the $\mu$-representation uniquely determines the in-degrees of all nodes, as long as we know which of the $\mu$-vectors belong to reticulations, and if we know, for each reticulation, the leaf for which it is stable (Proposition 2). A vertex $v$ is stable, if there exists a leaf $a$ such that all paths from the root to $a$ visit $v$.
We then ask if this can also be done for orchard networks. Because orchard networks do not contain tree-clones, it is easy to determine which $\mu$-vectors belong to reticulations in these networks, namely those which have multiplicity greater than 1. However, determining the stability of reticulations from the $\mu$-vectors is not obvious. In Section 4.2 we set out a way to do this. It turns out this can be done if one more condition is added: each reticulation must be the lowest reticulation above some bridge. At the end of Section 4.2.1, we show that if we know which $\mu$-vector belongs to the head of a bridge (which we call a bridge-node), then we can determine which $\mu$-vector belongs to the reticulation which is lowest above that bridge (Lemma 12). For these reticulations, it follows that we can identify the leaves with which they are stable. We finally give a characterization in Proposition 3 for the $\mu$-vectors of bridge-nodes in semi-binary stack-free networks. By combining these results with Proposition 2, it follows that for strongly reticulation-visible networks, the in-degrees are uniquely determined by the $\mu$-representation. Therefore, there is a bijection between the $\mu$-representation and the modified $\mu$-representation for these networks. It then follows from Theorem 1 what we set out to prove: strongly reticulation-visible orchard networks are encoded by their $\mu$-representation (Theorem 4).
@@ 4.1 Determining the in-degrees of reticulations
In this section we set out a relationship between the in-degrees of reticulations and the number of paths to the leaves below that reticulation. Naturally, higher in-degrees correspond to a larger number of paths crossing a reticulation. Therefore, it follows that the $\mu$-vectors of ancestors of a reticulation may contain information pertaining to the in-degree of that reticulation.
@@ 4.1.1 An equation relating the in-degrees of reticulations and $\mu$-vectors
Let $\mathcal{N}=(V,E)$ be an arbitrary X-DAG. Recall that $\rho$ denotes the root of the network.
@@
Let $a$ be an element of X. Let $R$ be the set of reticulations in $V$. Then,
$\mu(\rho)_{a}=\sum_{r_{i}\in R}(\delta^{-}(r_{i})-1)\mu(r_{i})_{a}+1.$
Note that $\mu(r_{i})_{a}=0$ if and only if $r_{i}$ is not an ancestor of $a$, which means that the contributions to the sum of 1 come from reticulation ancestors of $a$.
@@
Let us consider an $X$-DAG with $k=0$, i.e., a tree. If the hybridization number is zero then the graph contains no reticulations and therefore $R$ is empty. This also means there is a unique tree-path from the root to each leaf and thus $\mu(\rho)_{a}=1$. This shows the equation holds for each leaf of this graph.
Now suppose the equation holds for each leaf of any $X$-DAG with hybridization number lower than $k$. Let $\mathcal{N}$ be an $X$-DAG with reticulation set $R$ such that $h(\mathcal{N})=k$. Without loss of generality let $r_{j}$ be a highest reticulation in $\mathcal{N}$, which means $r_{j}$ has no reticulation ancestors. We can decrease the in-degree of $r_{j}$ by deleting an incoming edge $ur_{j}$. If $u$ is an elementary node we delete all edges on the maximal elementary path which visits $u$, as well as the nodes which become isolated by doing so. When we delete any of the incoming edges $ur_{j}$ in this way the resulting network $\mathcal{N}^{\prime}$ is still an $X$-DAG. Furthermore, the in-degrees of any other reticulations in the network have not decreased, which means $\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})=\delta^{-}_{\mathcal{N}}(r_{i})$ for any $i\neq j$. Moreover, the number of paths from any reticulation $r$ to any leaf $a$ has not changed and thus $\mu^{\prime}_{a}(r)=\mu(r)_{a}$. Finally, the hybridization number of $\mathcal{N}^{\prime}$ is equal to $k-1$ and by the induction hypothesis we have,
$\mu^{\prime}(\rho)_{a}=\sum_{r_{i}\in R^{\prime}}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu(r_{i})_{a}+1.$
@@
$\mu^{\prime}(\rho)_{a}$ $=\sum_{r_{i}\in R^{\prime}}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu(r_{i})_{a}+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu(r_{i})_{a}+(\delta^{-}_{\mathcal{N}}(r_{j})-2)\mu(r_{j})_{a}+1$
@@
$\mu^{\prime}(\rho)_{a}$ $=\sum_{r_{i}\in R^{\prime}}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu(r_{i})_{a}+1$
$=\sum_{r_{i}\in R}(\delta^{-}_{\mathcal{N}^{\prime}}(r_{i})-1)\mu(r_{i})_{a}+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu(r_{i})_{a}+(\delta^{-}_{\mathcal{N}^{\prime}}(r_{j})-1)\mu(r_{j})_{a}+1$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}_{\mathcal{N}}(r_{i})-1)\mu(r_{i})_{a}+(\delta^{-}_{\mathcal{N}}(r_{j})-2)\mu(r_{j})_{a}+1.$
@@
When we add the edge $ur_{j}$ (or the maximal elementary path which visits $u$ and ends in $r_{j}$) back to the network, we generate $\mathcal{N}$ from $\mathcal{N}^{\prime}$. In doing so we increase $\mu^{\prime}_{a}(\rho)$ by $\mu(r_{j})_{a}$ for any leaf $a$. Indeed, $r_{j}$ is chosen to be a highest reticulation in $\mathcal{N}$, and so there is a single path from the root to $r_{j}$ which uses the edge $(u,r_{j})$. By definition, there are $\mu(r_{j})_{a}$ paths from $r_{j}$ to any leaf $a$. This means that, for any leaf $a$
$\mu(\rho)_{a}$ $=\mu^{\prime}(\rho)_{a}+\mu(r_{j})_{a}$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu(r_{i})_{a}+(\delta^{-}(r_{j})_{\mathcal{N}}-2)\mu(r_{j})_{a}+1+\mu(r_{j})_{a}$
$=\sum_{r_{i}\in R\setminus\{r_{j}\}}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu(r_{i})_{a}+(\delta^{-}(r_{j})_{\mathcal{N}}-1)\mu(r_{j})_{a}+1$
$=\sum_{r_{i}\in R}(\delta^{-}(r_{i})_{\mathcal{N}}-1)\mu(r_{i})_{a}+1,$
@@
Note that this theorem holds for non-binary $X$-DAGs, as we never assumed a limit on the in-degree of reticulations or the out-degree of tree-nodes. 1 provides us with a system of linear equations that govern the in-degrees of reticulations as a function of $\mu$-vectors. We know this system of equations must have a solution as long as it belongs to a valid phylogenetic network. However, there are no guarantees yet that it has a unique solution. A network may contain more reticulations than there are linearly independent equations in the system. The network displayed in Figure 9 is semi-binary stack-free orchard, yet the system of equations generated by applying 1 to each leaf does not have a unique solution. Note that leaf $c$ has no reticulation ancestors and therefore the equation for leaf $c$ ($\mu(\rho)_{c}=1$) does not contribute. This means that there is one equation for leaf $a$ and one for leaf $b$, but three variables, the in-degrees of reticulations $r_{1},r_{2}$ and $r_{3}$. Note that the lower bound for the in-degree of reticulations is 2, and therefore, there are only a finite number of solutions. In this case there are two solutions to the system of equations and Figure 9 shows $(\delta^{-}(r_{1}),\delta^{-}(r_{2}),\delta^{-}(r_{3}))=(3,3,2)$ is the one belonging to this network.
Figure 9: A semi-binary stack-free orchard network for which the system of equations generated by applying 1 to each leaf does not have a unique solution. However, the solution $(\delta^{-}(r_{1}),\delta^{-}(r_{2}),\delta^{-}(r_{3}))=(3,3,2)$ is the unique one that belongs to this network.
@@ 4.1.2 In-degrees of stable reticulations
If we let $\mathcal{N}$ be a reticulation-visible X-DAG with reticulation set $R$ then the following is true. For $r\in R$, with leaf $a$ below $r$ such that all paths from the root to $a$ visit $r$, and $v$ any ancestor of $r$,
$\mu(v)_{a}=P_{vr}\mu(r)_{a},$ (1)
where $P_{vr}$ is the number of paths from $v$ to $r$. This is true, as every path from $v$ to $a$ must be the concatenation of a path from $v$ to $r$ and a path from $r$ to $a$. With this we gain the following lemma. This shows that in reticulation-visible networks, there is a unique equation for each reticulation.
@@
$\frac{\mu(\rho)_{a}}{\mu(r_{\ell})_{a}}$ $=\sum_{r\in A}(\delta^{-}(r)-1)\frac{\mu(r)_{a}}{\mu(r_{\ell})_{a}}+1$ (2)
@@
Let us generate $\mathcal{N}^{\prime}$ from $\mathcal{N}$ by first attaching a new leaf $a^{\prime}$ to $r_{\ell}$, by adding the edge $(r_{\ell},a^{\prime})$, and then adjusting the $\mu$-representation by adding a column for $a^{\prime}$, such that $\mu^{\prime}(v)=\mu(v)\oplus\mu(v)_{a^{\prime}}$. Note that this network is now an $X^{\prime}$-DAG, where $X^{\prime}=X\cup\{a^{\prime}\}$. Then by 1 we have:
@@
$\mu^{\prime}_{a^{\prime}}(v)=P_{vr_{\ell}}=\frac{\mu(v)_{a}}{\mu(r_{\ell})_{a}}.$ (4)
@@
$\frac{\mu(\rho)_{a}}{\mu(r_{\ell})_{a}}=\sum_{r\in A}(\delta^{-}(r)-1)\frac{\mu(r)_{a}}{\mu(r_{\ell})_{a}}+1.$
For the following lemmas, let $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ be two reticulation-visible X-DAG’s on the same leaf set $X$ with $\mu(\mathcal{N}_{1})=\mu(\mathcal{N}_{2})$. Let $R_{1}$ and $R_{2}$ denote the reticulation sets of $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, respectively. We say that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ agree on the reticulation set if there exists a bijection $f:R_{1}\rightarrow R_{2}$ such that $\mu(r_{1})=\mu(f(r_{1}))$ for all $r_{1}\in R_{1}$. For simplicity, we refer to $R_{1}$ (and by bijection, also $R_{2}$) as $R$. We assume that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ agree on the reticulation set. Finally, assume that for each reticulation $r\in R$, there exists a leaf $a\in X$ such that $r$ is stable with respect to $a$ in both networks.
@@
For any reticulation $r\in H$ the set $A$ in 4 contains only $r$ and therefore the in-degree of $r$ is given directly by Equation 2, applied to the leaf for which the reticulation is stable in both networks. ∎
@@
Let $H$ be the set of highest reticulations in $R$, which is the same set for $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$, by 5. By 6 the in-degrees of the reticulations in $H$ are the same in both networks. Let $S$ be the set of second highest reticulations in $R$, such that each reticulation $r\in S$ only has ancestors in $H$. By applying 7 the in-degrees of the reticulations in $S$ are also the same in both networks. Now, in the same way, the in-degrees of the third highest reticulations are equal and so on and so forth. This process must terminate as we only consider finite graphs. Thus, by repeated application of 7, we see that the in-degrees of all reticulations in $R$ are the same in both networks. ∎